OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − y = 6 4x − 2y = 12 a. 0 b. 4 c. x can be any number as there are infinitely many solutions to the system. d.There is no x value as there is no solution to the system.

OpenStudy (anonymous):

2x − y = 6 4x − 2y = 12 -----> divide by 2 2x − y = 6 2x − y = 6 These are the same equation, so both equation define the same line. Therefore there are infinitely many solutions to the system.